OpenStudy (anonymous):

represent ln5 as a definite intergral of the reciporcal function y=1/t

OpenStudy (anonymous):

\[\int\limits_{1}^{5} 1/t \]

OpenStudy (anonymous):

how would i grapth that to show the area it represents

OpenStudy (anonymous):

well you just need to graph 1/t and the area where x goes from 1 to 5

OpenStudy (anonymous):

how do i enter that in the calc

OpenStudy (anonymous):

with a graphic one you can do it. but I never used one

OpenStudy (anonymous):

do you know how 1/t looks like?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=graph+1%2Fx

OpenStudy (anonymous):

the area you need is the one under the graph from x=1 to x=5

OpenStudy (anonymous):

so the first one?

OpenStudy (anonymous):

practice makes perfect, if you do calculus you need to be able to draw basic graphs of functions

OpenStudy (anonymous):

so that is the the area? lol do i need to shade anything

OpenStudy (anonymous):

shade under the graph until the x axes

OpenStudy (anonymous):

perfect thanks

OpenStudy (anonymous):

what is the x-intercept of the tangent to y=e^x at x=45

OpenStudy (anonymous):

hmm, the derivative of e^x is e^x so the tangent at x=45 has slope of e^45 and the actual value is the same so then tangent line is y=e^45+e^45x the x intercept means that y=0 so 0=e^45+e^45x x=-1 that means that from x=45 you need to go back 1 to get the y intercept solution: x=44

OpenStudy (anonymous):

okay thats what i got. sweet

OpenStudy (anonymous):

well done!!